OpenStudy (anonymous):

what is the derivative of x^4(sec(x)-tan(x))? Help!!

OpenStudy (aum):

\[ \frac{d}{dx}x^4(\sec(x)-\tan(x)) = x^4~ *~ \frac{d}{dx}(\sec(x)-\tan(x)) ~+~ \frac{d}{dx}x^4 ~*~ (\sec(x)-\tan(x)) = ? \]

OpenStudy (aum):

\[ \frac{d}{dx}x^4 = ? \\ \frac{d}{dx}\sec(x) = ? \\ \frac{d}{dx}\tan(x) = ? \]Plug them above and simplify.